I'm confused...help?
If f(x) = x(x + 1)(x - 4), use interval notation to give all values of x where f(x)>0.

Answer by stanbon(75887)

If f(x) = x(x + 1)(x - 4), use interval notation
to give all values of x where f(x)>0.
--------
Find the values when f(x)=0
They are x=-1, or x=0, or x=4
Draw a number line and put the values -1,0,and 4 in proper order.
This determines 4 intervals; determines which intervals are part
of the solution set for f(x)>0 by choosing a test pt in each interval
and determining the sign of f(x)
-------
In (-inf,-1) choose x--100 where f(x)=-*-*- which is negative, so no solutions
In (-1,0) choose x=-1/2 where f(x)=-*+*- which is positive, so solutions here
In (0,4) choose x=2 where f(x)=+*+*- which is negative, so no solutions here
In (4,+inf) choose x=100 where f(x)=+*+*+ which is positive , so solutions here
---------
Therefore the solutions are in these intervals: (-1,0) union (4,+inf)
---------
Cheers,
Stan H.